OpenStudy (anonymous):

Which is the equation of the line with slope 5 that contains point (-2, -3)?

OpenStudy (anonymous):

equation of line -> 5x - y = 13

OpenStudy (galleara):

is it supposed to be in slope intercept form?

OpenStudy (anonymous):

ah, 5x - y +7 = 0

OpenStudy (galleara):

or it can also be put as y=5x+7

OpenStudy (anonymous):

yeah same..

OpenStudy (galleara):

yeah :)

OpenStudy (texaschic101):

y - y1 = m(x - x1)

OpenStudy (texaschic101):

are you familiar with this formula ?

OpenStudy (anonymous):

am.. and also used the same to get that answer.

OpenStudy (texaschic101):

good deal